# GIFT: Group-relative Implicit Fine Tuning Integrates GRPO with DPO and UNA

### v2

@@ 1 Introduction
Modern large language models (LLMs) are pretrained on trillions of tokens to acquire broad linguistic and factual knowledge [openai2024gpt4ocard]. Subsequently, supervised fine-tuning (SFT) further enhances their instruction-following capabilities [brown2020languagemodelsfewshotlearners]. Despite these advances, pretrained models often exhibit undesirable behaviors such as bias, unsafe outputs, and weaknesses in reasoning tasks like mathematics and programming.
To mitigate these issues, reinforcement learning (RL)–based alignment methods such as reinforcement learning from human feedback (RLHF) [ouyang2022traininglanguagemodelsfollow] and reinforcement learning with verifiable rewards (RLVR) [deepseekai2025deepseekr1incentivizingreasoningcapability] have been proposed. These approaches involve two stages: (1) training a reward model (RM) using the Bradley–Terry (BT) formulation on pairs of preferred and dispreferred responses, and (2) optimizing the policy with proximal policy optimization (PPO) [schulman2017proximalpolicyoptimizationalgorithms] to maximize cumulative reward while constraining divergence from the reference model. GRPO simplifies PPO by replacing the value function with the mean and standard deviation of reward values. While these online RL methods achieve strong alignment performance, they incur significant computational cost and training instability.
To simplify the alignment process, a parallel line of research has explored offline preference-based methods such as Direct Preference Optimization (DPO) [rafailov2024directpreferenceoptimizationlanguage]. DPO establishes a direct relationship between the implicit reward model and the optimal policy, reformulating RLHF into a single-stage, gradient-based optimization process that eliminates the need for online sampling and significantly reduces computational cost. UNA establishes a special relationship between implicit and explicit reward models, enabling heterogeneous supervision (e.g., preference, binary, or score-based data) through minimizing their discrepancy [wang2025unaunifyingalignmentsrlhfppo]. However, offline methods sacrifice exploration and adaptability, leading to weaker performance compared with online RL.
@@ 2.3 GIFT Derivation
For the explicit reward model, normalization can be performed following the GRPO workflow, yielding the normalized reward $r^{\prime}_{\phi}(x,y)$ as defined in Equation 2. Inspired by UNA, the LLM policy is then optimized by minimizing the discrepancy between the normalized implicit and explicit reward functions, as formulated in Equation 12. Although the temperature term $\beta$ is canceled during normalization, it should be reintroduced to balance exploration and exploitation in the RL objective (Equation 1). In practice, $\beta=1$ serves as a good default choice.
$L_{\text{GIFT-reward}}(\pi_{\theta})=\mathbb{E}_{(x,y)\sim D}\big[(r^{\prime}_{\phi}(x,y)-\beta\hat{r}^{\prime}_{\theta}(x,y))^{2}\big]$ (12)
@@ 3 Experiment
We first conduct RLVR-based comparisons between GIFT and GRPO using two backbone models: (1) deepseek-ai/deepseek-llm-7b-chat [deepseekai2025deepseekr1incentivizingreasoningcapability] and (2) Qwen/Qwen2.5-32B-Instruct [qwen2.5]. Experiments are performed on two benchmark datasets: (1) openai/gsm8k [cobbe2021gsm8k] and (2) DigitalLearningGmbH/MATH-lighteval [hendrycksmath2021].
Next, we extend the RLVR evaluation to a more challenging setting by training on the DAPO dataset [yu2025dapoopensourcellmreinforcement] and evaluating on the AIME dataset [aime2024_huggingface]. These experiments are conducted using Qwen/Qwen2.5-32B [qwen2.5] and Qwen/Qwen3-32B [yang2025qwen3technicalreport].
Finally, we compare GIFT and GRPO under the RLHF paradigm using the Infinity dataset [li2025infinityinstructscalinginstruction]. This evaluation is carried out with the same two models as in the initial RLVR experiments: deepseek-ai/deepseek-llm-7b-chat and Qwen/Qwen2.5-32B-Instruct. For reward model, Skywork/Skywork-Reward-V2-Llama-3.1-8B-40M is utilized [liu2025skywork].
Across most experiments, we keep the learning rate and the hyperparameter $\beta$ consistent to ensure fair comparisons. For GRPO, the learning rate is set to $1\times 10^{-6}$ with $\beta=0.001$. For GIFT, we use a learning rate of $3\times 10^{-6}$ and set $\beta=1$, noting that $\beta$ can be further tuned if necessary. However, in the RLVR experiments on the DAPO and AIME datasets with Qwen2.5-32B, training exhibited instability, and we therefore reduced the learning rates. Specifically, the learning rate is set to $3\times 10^{-7}$ for GRPO and $1\times 10^{-6}$ for GIFT in these settings.
For most cases, the learning rate and $\beta$ are the same. For GRPO, the learning rate is set to $1\times 10^{-6}$ and $\beta=0.001$. For GIFT, the learning rate is $3\times 10^{-6}$ and $\beta=1$, where $\beta$ can be further fine tuned if needed. However, in RLVR on DAPO and AIME dataset for Qwen2.5-32B, the training process is unstable and the corresponding learning rates are decreased. For GRPO, the learning rate is set to $3\times 10^{-7}$ while the learning rate for GIFT is set to $1\times 10^{-6}.$
@@ 3.3 Comparison between GIFT and GRPO in RLHF
Specifically, Table 2 reports results on seven diverse evaluation tasks: TruthfulQA [lin2022truthfulqameasuringmodelsmimic], BBQ [parrish-etal-2022-bbq], MBPP [austin2021programsynthesislargelanguage], ARC-Challenge [clark2018thinksolvedquestionanswering], Winogender [rudinger2018genderbiascoreferenceresolution], GPQA [rein2023gpqagraduatelevelgoogleproofqa], and MUSR [sprague2024musrtestinglimitschainofthought]. Across all seven tasks, GIFT consistently surpasses GRPO for both the 7B and 32B model sizes.
For generation-oriented evaluations such as AlpacaEval [dubois2024length] and Arena-Hard [arenahard2024], GIFT also outperforms GRPO in almost all settings. The only exception is the 7B style-control configuration, where GRPO achieves a marginally higher score than GIFT.
@@ A.1 RLHF and PPO
During reward model training, an explicit reward model is learned to predict a scalar score $r_{\phi}(x,y)$ for a given prompt–response pair $(x,y)$. The model is trained on pairwise preference data $(x,y_{w},y_{l})$, where $y_{w}$ denotes the preferred response and $y_{l}$ the dispreferred one. The probability that $y_{w}$ is favored over $y_{l}$, denoted as $P_{\phi}(y_{w}>y_{l}\mid x)$, is modeled using the Bradley–Terry (BT) framework, i.e., $P_{\phi}(y_{w}>y_{l}\mid x)=\frac{e^{r_{\phi}(x,y_{w})}}{e^{r_{\phi}(x,y_{w})}+e^{r_{\phi}(x,y_{l})}}=\sigma\!\left(r_{\phi}(x,y_{w})-r_{\phi}(x,y_{l})\right)$ where $\sigma(\cdot)$ denotes the logistic sigmoid function [bradley1952rank]. Intuitively, the higher the difference $r_{\phi}(x,y_{w})-r_{\phi}(x,y_{l})$, the greater the likelihood that the model predicts $y_{w}$ as more aligned with human preference. Eventually, the RM is trained through a binary cross entropy (BCE) loss $L_{\text{RM}}(\pi_{\theta})=-\mathbb{E}_{(x,y_{w},y_{l})\sim D}\left[\log\sigma\!\left(r_{\phi}(x,y_{w})-r_{\phi}(x,y_{l})\right)\right]$ [ouyang2022traininglanguagemodelsfollow].
@@
To stabilize training, Proximal Policy Optimization (PPO) [schulman2017proximalpolicyoptimizationalgorithms] is commonly adopted to optimize the objective in Equation 15. However, RLHF requires maintaining multiple large-scale components—namely, a policy model, reference model, reward model, and value model—resulting in substantial computational and memory overhead. These drawbacks significantly restrict the scalability and practicality of RLHF for LLMs.
@@ A.2 RLVR and GRPO
To reduce the computation cost of PPO, GRPO is then proposed [shao2024deepseekmathpushinglimitsmathematical, zheng2025groupsequencepolicyoptimization] as shown in Equation 16 where $r^{\prime}_{\phi}(x,y)$ refers to the normalized explicit reward model, $\mu_{\phi}$ refers to the mean of the explicit reward model and $\sigma_{\phi}$ refers to the standard deviation of the explicit reward model.
@@
When combined with the GRPO framework, as demonstrated in DeepSeek R1  [deepseekai2025deepseekr1incentivizingreasoningcapability], RLVR substantially enhances an LLM’s reasoning capacity. This integration enables the model to sustain longer and more coherent chains of thought (CoT), exhibit improved self-consistency across reasoning steps, and even display emergent “aha moments,” where the model identifies and corrects its own logical or computational errors during inference.
@@ A.3 DPO
To address the instability and inefficiency of RLHF and PPO, DPO was proposed as a simplified offline alternative utilizing the same reward model training’s dataset, i.e., $(x,y_{w},y_{l})$ in RLHF [rafailov2024directpreferenceoptimizationlanguage]. Based on the RL objective in Equation 15, DPO eliminates the explicit reward model and build a mapping between an implicit reward model and the optimal policy as shown in Equation 17
@@ A.4 UNA
One limitation in DPO is that it can only utilize pairwise dataset, while other format dataset cannot be utilized. In addition, the pairwise preference data provide less information compared with detailed information provided by reward mode. UNA extends DPO to pointwise dataset composed of (prompt, response, reward), i.e., $(x,y,r)$. In particular, UNA discovers that Equation 18 is also a special optimal implicit reward function for the RL objective in Equation 15 based on DPO’s implicit function in Equation 17 [wang2025unaunifyingalignmentsrlhfppo].

### v3

@@ 1 Introduction
Modern large language models (LLMs) are pretrained on trillions of tokens to acquire broad linguistic and factual knowledge [14]. Subsequently, supervised fine-tuning (SFT) further enhances their instruction-following capabilities [3]. Despite these advances, pretrained models often exhibit undesirable behaviors such as bias, unsafe outputs, and weaknesses in reasoning tasks like mathematics and programming.
To mitigate these issues, reinforcement learning (RL)–based alignment methods such as reinforcement learning from human feedback (RLHF) [15] and reinforcement learning with verifiable rewards (RLVR) [6] have been proposed. These approaches involve two stages: (1) training a reward model (RM) using the Bradley–Terry (BT) formulation on pairs of preferred and dispreferred responses, and (2) optimizing the policy with proximal policy optimization (PPO) [20] to maximize cumulative reward while constraining divergence from the reference model. GRPO simplifies PPO by replacing the value function with the mean and standard deviation of reward values. While these online RL methods achieve strong alignment performance, they incur significant computational cost and training instability.
To simplify the alignment process, a parallel line of research has explored offline preference-based methods such as Direct Preference Optimization (DPO) [17]. DPO establishes a direct relationship between the implicit reward model and the optimal policy, reformulating RLHF into a single-stage, gradient-based optimization process that eliminates the need for online sampling and significantly reduces computational cost. UNA establishes a special relationship between implicit and explicit reward models, enabling heterogeneous supervision (e.g., preference, binary, or score-based data) through minimizing their discrepancy [24]. However, offline methods sacrifice exploration and adaptability, leading to weaker performance compared with online RL.
@@
  • Eliminate hyperparameters requiring manual tuning—such as the clipping ratio and $\beta$.
@@ 2.3 GIFT Derivation
For the explicit reward model, normalization can be performed following the GRPO workflow, yielding the normalized reward $r^{\prime}_{\phi}(x,y)$ as defined in Equation 2. Inspired by UNA, the LLM policy is then optimized by minimizing the discrepancy between the normalized implicit and explicit reward functions, as formulated in Equation 12.
$L_{\text{GIFT-reward}}(\pi_{\theta})=\mathbb{E}_{(x,y)\sim D}\big[(r^{\prime}_{\phi}(x,y)-\hat{r}^{\prime}_{\theta}(x,y))^{2}\big]$ (12)
@@ 3 Experiment
We first conduct RLVR-based comparisons between GIFT and GRPO using two backbone models: (1) deepseek-ai/deepseek-llm-7b-chat [6] and (2) Qwen/Qwen2.5-32B-Instruct [23]. Experiments are performed on two benchmark datasets: (1) openai/gsm8k [5] and (2) DigitalLearningGmbH/MATH-lighteval [8].
Next, we extend the RLVR evaluation to a more challenging setting by training on the DAPO dataset [26] and evaluating on the AIME dataset [13]. These experiments are conducted using Qwen/Qwen2.5-32B [23] and Qwen/Qwen3-32B [25].
Finally, we compare GIFT and GRPO under the RLHF paradigm using the Infinity dataset [9]. This evaluation is carried out with the same two models as in the initial RLVR experiments: deepseek-ai/deepseek-llm-7b-chat and Qwen/Qwen2.5-32B-Instruct. For reward model, Skywork/Skywork-Reward-V2-Llama-3.1-8B-40M is utilized [12].
Across most experiments, we keep the learning rate consistent to ensure fair comparisons. For GRPO, the learning rate is set to $1\times 10^{-6}$ with $\beta=0.001$. For GIFT, we use a learning rate of $3\times 10^{-6}$ and set $\beta=1$. However, in the RLVR experiments on the DAPO and AIME datasets with Qwen2.5-32B, training exhibited instability, and we therefore reduced the learning rates. Specifically, the learning rate is set to $3\times 10^{-7}$ for GRPO and $1\times 10^{-6}$ for GIFT in these settings.
For most cases, the learning rate and $\beta$ are the same. For GRPO, the learning rate is set to $1\times 10^{-6}$ and $\beta=0.001$. For GIFT, the learning rate is $3\times 10^{-6}$ and $\beta=1$. However, in RLVR on DAPO and AIME dataset for Qwen2.5-32B, the training process is unstable and the corresponding learning rates are decreased. For GRPO, the learning rate is set to $3\times 10^{-7}$ while the learning rate for GIFT is set to $1\times 10^{-6}.$
@@ 3.3 Comparison between GIFT and GRPO in RLHF
Specifically, Table 2 reports results on seven diverse evaluation tasks: TruthfulQA [11], BBQ [16], MBPP [1], ARC-Challenge [4], Winogender [19], GPQA [18], and MUSR [22]. Across all seven tasks, GIFT consistently surpasses GRPO for both the 7B and 32B model sizes.
For generation-oriented evaluations such as AlpacaEval [7] and Arena-Hard [10], GIFT also outperforms GRPO in almost all settings. The only exception is the 7B style-control configuration, where GRPO achieves a marginally higher score than GIFT.
@@ A.1 RLHF and PPO
During reward model training, an explicit reward model is learned to predict a scalar score $r_{\phi}(x,y)$ for a given prompt–response pair $(x,y)$. The model is trained on pairwise preference data $(x,y_{w},y_{l})$, where $y_{w}$ denotes the preferred response and $y_{l}$ the dispreferred one. The probability that $y_{w}$ is favored over $y_{l}$, denoted as $P_{\phi}(y_{w}>y_{l}\mid x)$, is modeled using the Bradley–Terry (BT) framework, i.e., $P_{\phi}(y_{w}>y_{l}\mid x)=\frac{e^{r_{\phi}(x,y_{w})}}{e^{r_{\phi}(x,y_{w})}+e^{r_{\phi}(x,y_{l})}}=\sigma\!\left(r_{\phi}(x,y_{w})-r_{\phi}(x,y_{l})\right)$ where $\sigma(\cdot)$ denotes the logistic sigmoid function [2]. Intuitively, the higher the difference $r_{\phi}(x,y_{w})-r_{\phi}(x,y_{l})$, the greater the likelihood that the model predicts $y_{w}$ as more aligned with human preference. Eventually, the RM is trained through a binary cross entropy (BCE) loss $L_{\text{RM}}(\pi_{\theta})=-\mathbb{E}_{(x,y_{w},y_{l})\sim D}\left[\log\sigma\!\left(r_{\phi}(x,y_{w})-r_{\phi}(x,y_{l})\right)\right]$ [15].
@@
To stabilize training, Proximal Policy Optimization (PPO) [20] is commonly adopted to optimize the objective in Equation 15. However, RLHF requires maintaining multiple large-scale components—namely, a policy model, reference model, reward model, and value model—resulting in substantial computational and memory overhead. These drawbacks significantly restrict the scalability and practicality of RLHF for LLMs.
@@ A.2 RLVR and GRPO
To reduce the computation cost of PPO, GRPO is then proposed [21, 27] as shown in Equation 16 where $r^{\prime}_{\phi}(x,y)$ refers to the normalized explicit reward model, $\mu_{\phi}$ refers to the mean of the explicit reward model and $\sigma_{\phi}$ refers to the standard deviation of the explicit reward model.
@@
When combined with the GRPO framework, as demonstrated in DeepSeek R1  [6], RLVR substantially enhances an LLM’s reasoning capacity. This integration enables the model to sustain longer and more coherent chains of thought (CoT), exhibit improved self-consistency across reasoning steps, and even display emergent “aha moments,” where the model identifies and corrects its own logical or computational errors during inference.
@@ A.3 DPO
To address the instability and inefficiency of RLHF and PPO, DPO was proposed as a simplified offline alternative utilizing the same reward model training’s dataset, i.e., $(x,y_{w},y_{l})$ in RLHF [17]. Based on the RL objective in Equation 15, DPO eliminates the explicit reward model and build a mapping between an implicit reward model and the optimal policy as shown in Equation 17
@@ A.4 UNA
One limitation in DPO is that it can only utilize pairwise dataset, while other format dataset cannot be utilized. In addition, the pairwise preference data provide less information compared with detailed information provided by reward mode. UNA extends DPO to pointwise dataset composed of (prompt, response, reward), i.e., $(x,y,r)$. In particular, UNA discovers that Equation 18 is also a special optimal implicit reward function for the RL objective in Equation 15 based on DPO’s implicit function in Equation 17 [24].